Exercise 2-11

Pharoah Flower Cart incurred the following monthly costs.

Bouquets Sold
5,000 7,500 10,000
Balloons (20 per bouquet) $ 14,287 $ 16,709 $ 19,134
Insurance $ 5,000 $ 5,000 $ 5,000
Delivery $ 5,782 $ 7,934 $ 10,084
Employee compensation $ 13,513 $ 14,723 $ 15,948
Advertising $ 1,500 $ 1,500 $ 1,500


Using the form y = mx + b, estimate the cost formula for each cost incurred by Pharoah Flower Cart. (Round unit cost to 2 decimal places, e.g. 52.75. If the cost equation has no variable component enter 0 for the amount.)

Balloons y = $

x
+ $

Insurance y = $

x
+ $

Delivery y = $

x
+ $

Employee compensation y = $

x
+ $

Advertising y =    $

x
+ $
Answer #1

Balloons Variable Cost = 16709 -14287 7500 - 5000 2422 = 0,9688 2500 fixed cost = 16709- 7500 X 0.9688 = 9443 I y = 0.9688 X